OpenStudy (anonymous):

how do i find the derivative of this using the chain rule please help somebody\[F(x)=x \sqrt{x ^{2}+1}\]

OpenStudy (bahrom7893):

Okay first throw the x into the sqrt.

OpenStudy (bahrom7893):

F(x) = sqrt(x^4 + x^2)

OpenStudy (bahrom7893):

Or F(x) = (x^4+x^2)^(1/2)

OpenStudy (bahrom7893):

Now take the derivative: F'(x) = (1/2) * (x^4 + x^2 )^(-1/2) * (4x^3 + 2x)

OpenStudy (bahrom7893):

\[f'(x) = (4x^3+2x)/(2\sqrt(x^4+x^2))\]

OpenStudy (anonymous):

can u please explain the part when u threw the x into the square root

OpenStudy (bahrom7893):

okay x is the same as \[\sqrt{x^2}\]

OpenStudy (anonymous):

yep

OpenStudy (bahrom7893):

so to put it inside the sqrt, multiply everything inside by x^2

OpenStudy (anonymous):

ok great thank u soo much appreciate the help :)

OpenStudy (bahrom7893):

Can you click on Become a fan next to my name, please! Thanks =)

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

can i ask u another question lol ?